One part of 10.000 investment at rate 7.5%/year. The other part at rate 8.5% per year. Income from the two investments = 820. How much is the investment amount for each rate?
Answer by Boreal(15235)

x at 0.075
10,000-x at 0.085
They add to 820
.075x+0.085(10,000-x)=820
.075x+850-0.085x=820
-0.01x=-30
multiply both sides by -100
x=$3000@0.075=$225
10,000-x=$7000@0.085=$595
They add to $820.